### v1

# Unveiling Challenges for LLMs in Enterprise Data Engineering

Abstract.
Large Language Models (LLMs) have demonstrated significant potential for automating data engineering tasks on tabular data, giving enterprises a valuable opportunity to reduce the high costs associated with manual data handling. However, the enterprise domain introduces unique challenges that existing LLM-based approaches for data engineering often overlook, such as large table sizes, more complex tasks, and the need for internal knowledge. To bridge these gaps, we identify key enterprise-specific challenges related to data, tasks, and background knowledge and conduct a comprehensive study of their impact on recent LLMs for data engineering. Our analysis reveals that LLMs face substantial limitations in real-world enterprise scenarios, resulting in significant accuracy drops. Our findings contribute to a systematic understanding of LLMs for enterprise data engineering to support their adoption in industry.
The source code, data, and/or other artifacts have been made available at https://github.com/DataManagementLab/llmeval-enterprise-challenges.
1. Introduction
Data engineering has high manual overheads.
Large enterprises generate vast amounts of tabular data that drives applications like machine learning and analytical query processing. Data engineering is crucial for understanding this raw data and transforming it into a suitable form for its downstream usage. It encompasses a range of tasks, from data exploration and transformation to data integration and cleaning. Since such tasks often impose significant manual overhead to apply existing tools to the specific data at hand, the automation of individual data engineering tasks like entity matching (Konda et al., 2016; Mudgal et al., 2018) and column type annotation (Hulsebos et al., 2019; Zhang et al., 2020a) with the help of machine learning has long drawn attention from researchers. Nevertheless, adapting such machine learning approaches to new datasets and tasks often requires computer science expertise, rendering them inaccessible to many practitioners.
LLMs to the rescue?
Recent work has shown that Large Language Models (LLMs) such as GPT-4 (OpenAI et al., 2024) can be directly applied to data engineering tasks on tabular data, indicating that they achieve state-of-the-art results on various table-based tasks without requiring task-specific architectures and training (Narayan et al., 2022; Jaimovitch-López et al., 2022; Kayali et al., 2024; Barke et al., 2024). Their out-of-the-box nature is a significant advantage over other machine learning approaches that require supervised training for each dataset and task. One example relevant for data understanding is the task of column type annotation, where the goal is to annotate the columns of a relational table with semantic types from a given ontology. Whereas machine learning-based approaches like Sherlock (Hulsebos et al., 2019) and Sato (Zhang et al., 2020a) require re-training for each new set of semantic types, LLMs can easily support different sets of semantic types by including them in the prompt (Korini and Bizer, 2023). Therefore, they could provide a promising avenue to automate data engineering tasks.
Is this also true for enterprise data?
Recently, some first papers have made the point that while LLMs achieve good results on existing benchmarks, they fail miserably when applied to real-world enterprise data (Kayali et al., 2025; Bodensohn et al., 2024b). A central observation here is that the data in public benchmarks is often crawled from web resources like Wikipedia (Bhagavatula et al., 2015) and GitHub (Hulsebos et al., 2023). In contrast, the data from companies running their business processes with software systems like those from SAP differs fundamentally from these datasets in many aspects, including table sizes, sparsity, and data types (Kayali et al., 2025; Savelieva et al., 2022; Vogelsgesang et al., 2018). Since LLMs are typically trained on public data scraped from the web (Brown et al., 2020; Meta, 2024), they have not seen significant amounts of such enterprise data during their training.
A significant performance decline on enterprise data.
In this paper, we thus set out the goal of systematically studying the performance of LLMs for enterprise data engineering. As a “sneak peak” into our results, we demonstrate the results of a first experiment on enterprise data where we compare the performance of LLMs on the task of column type annotation on a public benchmark to a dataset of real-world customer data from SAP. As shown in Figure 1 (left bar group), LLMs of different types and generations achieve high F1 scores of up to 0.94 when solving this task on a public corpus like the SportsTables dataset (Langenecker et al., 2023). In contrast, as shown in Figure 1 (second bar group), using LLMs on the representative customer data from SAP for column type annotation with a comparable number of semantic types causes a substantial performance decrease across all models, ranging from single-digit F1 scores of 0.07 up to 0.38.
Beyond data challenges.
Beyond these data challenges, data engineering in enterprises faces additional difficulties:
(1) Task complexity: Enterprise tasks are often more complex than their academic formulations (Savelieva et al., 2022). For example, literature about the task of entity matching often assumes that each entity is one row in a single table (Konda et al., 2016; Peeters and Bizer, 2023). By contrast, entities in enterprises are often business objects that span across multiple tables, making matching notoriously more difficult. Moreover, the tasks themselves are often more complex. For example, increasing the task complexity of column type annotation by scaling to the true number of semantic types (5,089 instead of 200) in the SAP system leads to a further 50% drop in F1 scores, as shown in Figure 1 (+ Task Challenges).
(2) Internal knowledge: Data engineering in enterprises also often requires internal knowledge that is absent from public sources, limiting LLMs’ capabilities to understand the data without additional knowledge. This is especially true for schema customizations, which involve customer-defined semantic types. On SAP’s column type annotation dataset, when adding those columns to the task, F1 scores for such customer-defined columns are near zero, as shown in Figure 1 (+ Knowledge Challenges).
A broad analysis of enterprise data engineering.
In this paper, we conduct a broad experimental analysis on representative customer data and case studies reflecting real enterprise scenarios. Our goal is to unmask the jagged out-of-the-box capabilities of LLMs to provide insights into where they can be reliably used and where additional tweaks and improvements are necessary. To accomplish this, we apply five LLMs from three model providers to various established data engineering tasks. As the main goal of this paper, we want to highlight how particular enterprise-specific challenges impact LLM performance and provide insights that can guide future efforts to make LLMs viable for enterprise data engineering.
We believe this to be the first attempt to examine LLMs for data engineering on real enterprise data at this breadth. We see it as an important first step (of many) to make LLMs viable for enterprise data engineering.
Contributions.
Our key contributions in this paper are:111This paper extends our previously published work (Bodensohn et al., 2024b, a) on this topic. (1) We systematically analyze the challenges involved in enterprise data engineering and structure them into aspects regarding enterprise data, enterprise tasks, and enterprise knowledge. (2) We experiment on representative enterprise data to show how it differs from existing public benchmarks and understand its impact on LLM performance. (3) We conduct multiple case studies that reflect real-world enterprise scenarios, allowing us to extensively evaluate how different challenges affect LLM performance in isolation. (4) We discuss potential directions for addressing these challenges, as well as the costs of using LLMs at enterprise scale. (5) Our experiments are performed with five recent LLMs from three model providers (OpenAI, Anthropic, and Meta). To enable follow-up research, we make the code—including our full evaluation setup, all prompts, and where legally possible also our data—of this paper available to the broader research community.
Outline.
The paper is structured as follows: Section 2 discusses the setting of our study on LLMs for data engineering, and Section 3 introduces our methodology for analyzing the challenges of using LLMs in enterprises. Sections 4 through 6 then present the analysis, experimental results, and discussions for the identified categories of challenges: data, tasks, and knowledge. Section 7 discusses the costs involved in enterprise-scale data engineering. Finally, Section 8 summarizes our conclusions and avenues for future research.
| Context Window | Reason- ing | USD Per 1M | ||
| Input | Output | |||
| GPT-4o-Mini (2024-07-18)1 | 128K | no | ||
| GPT-4o (2024-08-06)1 | 128K | no | ||
| o1 (2024-12-17)1 | 200K | yes | ||
| Claude 3.5 Sonnet (v2)2 | 200K | no | ||
| Llama 3.1 Instruct (70B)3 | 128K | no | ||
Pricing by OpenAI, Anthropic, and AWS Bedrock3 in Feb 2025.
2. Setting of our Study
Using LLMs to solve table-based tasks is a promising research direction that has been actively studied in recent years (Narayan et al., 2022; Zhang et al., 2024a; Peeters and Bizer, 2023; Korini and Bizer, 2023; Zhang et al., 2024b). This section introduces the models we evaluate and briefly summarizes existing research on LLMs for data engineering.
Choosing LLMs for the study.
LLMs are text foundation models trained on large corpora to complete natural language inputs (Brown et al., 2020) and follow user instructions (Wei et al., 2022; Liu et al., 2023). Modern LLMs like GPT-4 (OpenAI et al., 2024) and Llama 3 (Grattafiori et al., 2024) build on a range of further ideas, from mixture-of-experts architectures (Shazeer et al., 2017) to more efficient implementations (Dao et al., 2022), leading to continuous improvements in abilities like applying background knowledge and handling long inputs. With the release of reasoning models such as OpenAI’s o1 (OpenAI, 2024b) and Deepseek’s R1 (DeepSeek-AI et al., 2025), LLMs perform even better at tasks that require several intermediate steps. While many LLMs would be relevant candidates to include in our evaluation, our selection is necessarily limited. To ensure generalizable results, we evaluate five recent LLMs from three model providers covering multiple sizes and costs and including open and closed models. As shown in Table 1, we use GPT-4o (OpenAI, 2024a) as the state-of-the-art and GPT-4o-Mini as the cheapest model from OpenAI as well as o1 (OpenAI, 2024c) as a reasoning model. We further include Claude 3.5 Sonnet (Anthropic, 2024) from Anthropic and Llama 3.1 Instruct (Meta, 2024) from Meta in our evaluations.
Positioning this study.
Although several papers have already observed the differences between enterprise and web data (Vogelsgesang et al., 2018; Savelieva et al., 2022; Kang and Naughton, 2003; Zhu et al., 2019), existing research on LLMs for data engineering primarily uses evaluation datasets based on tables from public web sources, calling the applicability of LLMs on real-world enterprise data into question. More recently, some enterprise-specific benchmarks have been released for tabular prediction tasks (Klein et al., 2024), Text-to-SQL (Chen et al., 2024), and column type annotation (Kayali et al., 2025). Related to our work, in a recent study, Kayali et al. (2025) quantify performance gaps between private and public data for the task of semantic column type annotation and find that benchmarks based on public data overestimate the performance of LLMs. However, these studies focus only on individual tasks and consider the data as the only difference in enterprise scenarios. In contrast, we perform case studies on various diverse tasks to systematically analyze the challenges along several dimensions (data, tasks, and knowledge) of working in enterprise scenarios.
Working with real enterprise data.
Attempts like ours often fail because enterprise data is usually highly confidential and, therefore, hard to use in evaluations. For this paper, we were able to experiment with actual customer data from the enterprise systems of SAP. While one could argue that this study is still limited because we only use data from SAP, and there are clearly many alternative enterprise systems, SAP stands out as a dominant player in enterprise software systems across multiple industries worldwide. As such, we believe that our insights based on SAP data are highly valuable on their own and hope that our paper inspires other researchers with access to similar enterprise datasets to repeat our evaluations on their data. Importantly, the core characteristics of SAP data reflect the findings from other papers observing the differences between enterprise and web data (Vogelsgesang et al., 2018; Savelieva et al., 2022; Kang and Naughton, 2003; Zhu et al., 2019), highlighting the generality of our results beyond SAP data.
3. Design of our Study
| Tables | Columns | Rows | Sparsity1 | Data Types2 | Column Type Annotation | |||||
| Med | 95th | Med | 95th | abc | 123 | Column Types | Labeled Columns | |||
| WikiTables-TURL | 397,098 | 1 | 3 | 8 | 43 | 0.12 | 1.00 | 0.00 | 255 | 628,254 |
| SOTAB | 59,548 | 7 | 17 | 33 | 721 | 0.08 | 0.85 | 0.15 | 91 | 162,351 |
| GitTablesCTA | 1,100 | 12 | 33 | 25 | 263 | 0.12 | 0.33 | 0.67 | 122 — 593 | 2,517 — 1,3743 |
| SportsTables | 1,183 | 21 | 31 | 32 | 924 | 0.07 | 0.16 | 0.84 | 452 | 24,821 |
| SAPCTA | 1004 | 46 | 343 | 473,038 | 50,836,964 | 0.43 | 0.55 | 0.455 | 5,089 | 8,106 |
1
Sparsity is the fraction of empty cells.
2
Non-numeric (abc) and numeric (123) columns determined by pandas.
3
Using semantic types from DBPedia — Schema.org.
4
We experiment on a representative sample from the thousands of tables in the customer system.
5
Only 14% of columns have numerical SQL types like INT and DECIMAL.
We aim to take a holistic view of data engineering in enterprises by systematically analyzing the out-of-the-box performance of LLMs, covering enterprise-specific challenges along the dimensions of data, tasks, and knowledge. Figure 2 provides an overview of the tasks we have chosen for analyzing the performance along each dimension. Below, we explain the rationale for selecting these tasks as examples to examine challenges that arise broadly across different enterprise tasks and scenarios. Sections 4 to 6 then describe our experimental evaluation for each of these challenges.
The data challenge.
Enterprise data differs from public evaluation datasets in various aspects. First, tables in enterprise databases are substantially larger in their number of columns and rows, and the schema and data are often more complex with table names, column names, and values that lack intuitive meaning. Moreover, enterprise data exhibits a much higher sparsity compared to public data, leaving many cells empty. Together, these factors make data engineering on enterprise data much more challenging. To explore these challenges, in our study, we focus on the task of column type annotation (see Figure 2 left) (Hulsebos et al., 2019; Zhang et al., 2020a). We have chosen this task because it is well-studied in the literature and thus enables the comparison of accuracies with public datasets. Moreover, the task can be formulated in various alternative ways, stressing different data factors (e.g., with and without schema information as input), allowing us to study the effects of different challenges like table size, sparsity, and descriptiveness on the data and schema level.
The task challenge.
Beyond the complexities of enterprise data, enterprise tasks themselves are also more complex. First, while academia typically studies data engineering tasks in isolation, enterprise tasks in practice are often compounds of multiple simpler data engineering steps. As such, we argue that the evaluation procedures themselves must change and we need to study accuracy end-to-end and analyze effects such as how errors propagate. In this paper, we examine the compound nature of enterprise tasks using the example of integrating two customer databases, as illustrated in Figure 2 (middle). This process involves the tasks of schema matching, entity matching, and data integration. A second challenge for enterprise tasks (not shown in Figure 2) is that even the individual steps are often more complex. For example, while entity matching in academia assumes 1:1 matches across rows of two tables (Peeters et al., 2024; Mudgal et al., 2018), matching in enterprise scenarios covers more complex scenarios like mapping several bank transfers to one invoice, and the required data is often scattered across multiple tables.
The knowledge challenge.
As a last challenge, data engineering in enterprises often requires enterprise-specific knowledge. This is particularly challenging because LLMs will likely not have seen the required information during training since it is covered only in internal documentation or even just certain implementation details in the code of enterprise systems. To analyze such challenges, we have selected a task related to data exploration, which requires translating natural language queries about business processes into the enterprise-specific query language SIGNAL, which differs slightly from SQL (Figure 2 right) (Kampik et al., 2023). The task is interesting because LLMs should, judging by its similarity to Text-to-SQL, in principle be able to solve it. However, they lack particular information about how exactly SIGNAL differs from SQL. At the core, the question we want to answer here is to what extent the lack of enterprise knowledge affects the accuracy of LLMs, and whether it can be increased by providing enterprise-specific knowledge as context to LLMs.
An orthogonal challenge: cost.
Since enterprise tables are large and can contain millions of rows, using LLMs on such data to solve complex multi-step tasks can cause high costs, rendering some of the larger, more complex LLMs economically unviable. We discuss this aspect in Section 7.
4. The Data Challenge
In this section, we quantify the anatomy of enterprise data by comparing real-world data from SAP databases to publicly available table corpora. We point out four challenges (C1-C4) specific to enterprise data and perform experiments to evaluate how they affect LLM performance. As an example task, we choose the task of column type annotation shown in Figure 2 (left).
4.1. Data Challenges
For this study, we constructed a new corpus SAPCTA . Table 2 compares the data characteristics of the real-world customer data in SAPCTA to several publicly available column type annotation benchmarks — WikiTables-TURL (Deng et al., 2020), SOTAB (Korini et al., 2022), GitTablesCTA (Hulsebos et al., 2021), and SportsTables (Langenecker et al., 2023). We observe the following differences:
C1: Table size. A first important observation is that enterprise tables typically have substantially more rows and columns than the tables in public corpora. As shown in Table 2, some tables have hundreds of columns and millions of rows. While the large scale is a well-established data management problem (Zhu et al., 2019), it poses challenges for LLMs, which have limited context windows. Although recent models have extended context windows, feeding large tables into LLMs still has downsides since latency and cost depend on the input size, and recent studies have shown that long contexts can lead to degraded performance for data residing in the “middle” (Liu et al., 2024a).
C2: Descriptiveness. Another important insight is that schema properties like table and column names are often not descriptive but rather abbreviations that can only be understood with background knowledge or additional metadata (Kang and Naughton, 2003). This additional metadata is often unavailable or may not fit into the context window of the LLM. Moreover, the background knowledge is often specific to the particular enterprise, causing challenges for LLMs trained exclusively on publicly available data, as we discuss in Section 6.
C3: Sparsity. A third insight is that enterprise data is highly sparse. Table 2 shows that on average, 43% of the cells in enterprise tables are empty, compared to only 7-12% in existing datasets. This high level of sparsity results in a significant lack of information, which poses a challenge for LLMs that rely on contextual cues to make accurate predictions. Moreover, we find that in addition to empty values, the cells in enterprise tables often contain dummy values such as 00000 that also denote the absence of an actual value.
C4: Data types. Surprisingly, we find that only 45% of the columns in SAPCTA are classified as numerical data by pandas, challenging the common assumption that enterprise data is predominantly numerical (Langenecker et al., 2023). Moreover, we see that only 14% of columns in the database schema have numerical data types like INT and DECIMAL. A closer inspection of the actual data reveals that the non-numerical data type NVARCHAR is often used to store symbolic values and codes such as invoice and material numbers, which is in line with previous findings (Vogelsgesang et al., 2018). Since these values are not self-expressive, LLMs cannot make use of them without additional context.
4.2. Experiments & Results
To study how the challenges C1-C4 affect LLMs for data engineering, we compare the model performance on our real-world enterprise dataset SAPCTA to the performance on existing evaluation datasets.
Column type annotation task.
For the evaluation, we have chosen the task of column type annotation, as it is a well-established task where the goal is to annotate the columns of a relational table with semantic types from a pre-defined ontology (Hulsebos et al., 2019; Zhang et al., 2020a). We see it as an interesting example task to uncover the challenges of understanding enterprise tables with LLMs because it requires a semantic understanding of the content of each column as well as the values of other columns and the table and column names, which can all provide important signals to derive a semantic type.
Setup.
Our SAPCTA corpus spans diverse business domains such as Finance, Sales and Distribution, Material Management, and Production Planning. For our experiments, we select 100 representative tables from the larger corpus, which contains multiple thousands of tables. Our prompting strategy builds on best practices from existing literature, where the model annotates the columns of a given table based on a list of semantic types in the prompt and one randomly selected example (Zhang et al., 2020a; Korini and Bizer, 2023). Due to the LLMs’ limited context windows, we have to limit each table to three randomly selected rows. For similar reasons, we serialize the tables in CSV format, which requires fewer formatting tokens than other serialization schemes like Markdown and JSON (Singha et al., 2023; Sui et al., 2024). We instruct the model to generate the column types as a JSON-formatted list.
Exp. 1: Enterprise vs. public tables. In our first experiment, we compare the performance for column type annotation on our SAPCTA corpus with the performance on GitTablesCTA (Hulsebos et al., 2021) and SportsTables (Langenecker et al., 2023). We perform each experiment twice, with and without including the table and column names (i.e., the table schema) in the prompt. Existing evaluations typically leave out the column names, since the semantic types are directly derived from them and the task would thus become trivial. However, for SAPCTA , the task is much harder. Therefore, we want to investigate how much the additional information helps.
Table 3 shows the results of this experiment. We make the following key observations: (1) LLMs have severe issues with column type annotation on enterprise data, leading to substantially worse results compared to the web resources GitTablesCTA and SportsTables. Especially in the experiments without table and column names, the results on enterprise data are particularly poor (F1 scores of only up to 0.05), indicating that the enterprise data on its own contains few helpful signals. (2) Adding table and column names to the prompt improves the results on the enterprise data, but only up to 0.34 using Claude 3.5 Sonnet, which is still much lower than for web tables. The remaining performance gap could potentially be attributed to the non-descriptive schema, the extremely wide and sparse tables, and the complex data types described in Section 4.1.
| GitTablesCTA | SportsTables | SAPCTA | ||||
| Column Names | w/out | with | w/out | with | w/out | with |
| GPT-4o-Mini | ||||||
| GPT-4o | 0.99 | |||||
| Claude 3.5 Sonnet | 0.67 | 0.68 | 0.94 | 0.05 | 0.34 | |
| Llama 3.1 Instruct | ||||||
Exp. 2: Textual vs. numerical data. LLMs are known to often perform better on textual data than on numerical data (Frieder et al., 2023; Langenecker et al., 2023). To analyze this effect on enterprise data, we examine the performance for non-numerical and numerical columns in our SAPCTA corpus. Table 4 shows that, as expected, we see a higher performance on non-numerical enterprise data. By contrast, the results on the public benchmarks GitTablesCTA and SportsTables are inconclusive, with only small performance gaps between non-numerical and numerical columns that differ between models. The low performance on the SAPCTA dataset (F1 scores of only up to 0.41 for non-numerical and 0.27 for numerical data) indicates that numerical data in enterprise systems is even harder to understand than in web tables. Furthermore, the low scores for non-numerical columns stems from the fact that the enterprise tables often store identifiers like INV0014056 as type NVARCHAR, as explained in Section 4.1 (C4).
| GitTablesCTA | SportsTables | SAPCTA | ||||
| Data Types | abc | 123 | abc | 123 | abc | 123 |
| GPT-4o-Mini | ||||||
| GPT-4o | 0.99 | 0.98 | 0.87 | |||
| Claude 3.5 Sonnet | 0.97 | 0.41 | 0.27 | |||
| Llama 3.1 Instruct | ||||||
Exp. 3: Table width and sparsity. To further investigate the performance gap between web tables and enterprise data, we incrementally adapt the enterprise tables to resemble the characteristics of web tables more closely. Since two of the main differences are table width and sparsity (see C1 and C3), we vary the number of columns per table by randomly sampling subsets of columns and vary the sparsity by initially selecting only non-sparse columns and randomly removing individual cell values. Figure 3 shows that increasing numbers of columns lead to substantially worse results, indicating that the large table widths in enterprise data are indeed a major problem for LLMs. Regarding sparsity, Figure 4 shows worse results with increased sparsity if no table and column names are provided, whereas with table and column names, increased sparsity does not change the results much. This indicates that for the enterprise data, LLMs rely primarily on the column headers to predict the semantic types and do not take the cell values into account.
4.3. Discussion
Learnings.
Our analysis of the representative customer data from SAP shows that enterprise data has fundamentally different characteristics, and our experiments demonstrate that these differences lead to substantial performance declines. We have seen that the performance decreases with higher sparsities and growing table sizes, indicating that LLMs are unreliable at scale and sometimes fail to uphold even the basic table structure by generating an incorrect number of labels. While we have shown this for the task of column type annotation, these data challenges will appear again throughout the following sections for other data engineering tasks like schema matching and entity matching, highlighting that they are, in fact, general and affect a broad range of tasks.
Outlook.
We believe that overcoming these challenges requires better representations for enterprise data. A first promising direction is the development of foundation models for relational data (Vogel et al., 2022; Wu et al., 2025). While there is already a large body of research on foundation models for tabular data (Qu et al., 2025; He et al., 2025; Li et al., 2024; Yang et al., 2022; Eisenschlos et al., 2021; Herzig et al., 2020; Yin et al., 2020; Deng et al., 2020), computing representations for complex enterprise data is still an open research problem. Challenges include topics such as scalability and cost as well as complex data types and non-descriptive table and column names.
To mitigate technical problems such as context size limitations, we could for example apply windowing approaches to handle the large tables. However, this approach is clearly limited as important information may be lost, which is especially problematic due to the low self-expressiveness of the enterprise data. One way to alleviate this issue could be to better contextualize the raw data by integrating metadata such as data dictionaries, which contain textual descriptions for table and column names and for symbolic values. However, preparing such additional data for LLMs currently imposes high manual overheads, as we will discuss in Section 5.
5. The Task Challenge
A second dimension where data engineering is different in enterprise scenarios is the complexities of the tasks, as shown in Figure 2 (center). In particular, the academic definition of a task often makes several simplifying assumptions that do not hold in real-world scenarios. For example, when looking at entity matching, the assumption is typically made that we compare the similarity of individual rows of two tables (Peeters and Bizer, 2023; Konda et al., 2016). However, when looking at entity matching in enterprises, business entities often span across multiple tables, forming a graph of tuples. Moreover, enterprises typically approach data engineering with broad business goals in mind. Therefore, enterprise tasks are often composed of multiple individual steps and are thus often compounds of simpler tasks (e.g., combining schema matching and entity matching). In the following, we present a detailed discussion of how enterprise tasks differ in their nature from the complexity of tasks evaluated in academia.
5.1. Task Challenges
We base our analysis of task-specific challenges in enterprise data engineering (C5-C8) on the following two case studies.
Case study 1: enterprise entity matching.
Entity matching is a task that often occurs in enterprise scenarios. In the following, we study a representative scenario of matching the bank statements of incoming payments to open invoices, which is shown in Figure 5. The scenario showcases many challenges in enterprise entity matching, as discussed below, including problems such as complex table structures and the fact that matching requires finding 1:N or even N:M matches.
Case study 2: enterprise database integration.
For our second case study, we look into another enterprise scenario where two databases from different companies must be integrated. To ensure a unified and consistent customer database, this process involves the individual steps of schema matching, customer record matching, and data integration, as illustrated in Figure 2 (center).
In the following, we highlight the general challenges that arise in enterprise tasks and provide examples based on these two scenarios.
C5: Entities span multiple tables. As mentioned before, entities in enterprise systems are often business objects that are represented by multiple rows stored in different connected tables. For example, in SAP systems, data pertaining to a particular material is scattered across the MARA (material type and basic statistics), MARC (manufacturing-related details), MBEW (valuation data), and other tables. The same holds true for our first case study (entity matching), which we later show in our evaluation. As such, for many data engineering tasks, including entity matching, one must either manually construct views that extract the fields relevant to the task into a single table, or approaches have to deal directly with the complex table structures that form a business entity.
C6: Compound tasks. While data engineering tasks in research are usually addressed as isolated problems, such as deduplication (Papenbrock et al., 2015) and missing value imputation (Mei et al., 2021), tasks in enterprise contexts are typically approached on a more holistic level concerning broader business objectives. Instead of focusing on individual tasks, enterprises aim to solve end-to-end workflows, as highlighted by our second case study (data integration). While steps of end-to-end workflows can be executed sequentially, errors often propagate and amplify in later steps. As such, analyzing steps in isolation does not reveal the overall quality of the task in enterprise scenarios.
C7: Task complexity. Moreover, even individual workflow steps can be more complex than their counterparts in the academic setting. In the following, we discuss these differences by the example of entity matching: (1) Matching between different types of entities: In public entity matching datasets, the entities to be matched are usually of the same type, such as e-commerce products, restaurants, and scholarly articles (Konda et al., 2016; Peeters et al., 2024). Meanwhile, enterprise scenarios often require matching between different types of entities, like products to commodity codes or payments to invoices (as in our first case study), which have overlapping but different sets of attributes, as shown in Figure 5. (2) Multi-matches: An additional challenge in enterprise scenarios is that the matches are often not 1:1 matches as in the literature but can also be 1:N, N:1, or even N:M matches, making the problem much more difficult. In our payment-to-invoice matching scenario, it is quite common for a customer to pay multiple invoices with only one payment (1:N), or for one invoice to be paid by multiple payments (N:1), such as in the case of down payments or by holding back money due to quality issues with the product. Even a combination of both cases is possible (N:M).
C8: Data complexity on top. Along with these task complexities, the complexity of the data makes the tasks themselves also harder. For example, when matching payments to invoices (first case study), the memo lines of incoming bank statements are not standardized, but rather free-form text entered by customers. Therefore, complex errors occur regularly, making automated matching difficult. Compared to such data errors in enterprise scenarios, the data errors in academic entity matching datasets are often much simpler.
5.2. Experiments & Results (Case Study 1)
To empirically examine the challenges that arise when solving enterprise data engineering tasks with LLMs, we first conduct experiments for our first case study with the payment-to-invoice matching scenario, which allows us to show many of the issues emerging from challenges C5, C7, and C8 discussed before. Experiments for the second case study, which focuses on challenge C6 (compound tasks), follow in the next subsection. For our experiments, we are using a dataset based on real-world customer data. Moreover, we pre-processed the dataset so that we can systematically include the challenges discussed before and analyze the performance when incrementally adding these challenges (e.g., with and without 1:N matches, with and without data errors).
Setup.
For the analysis, we use a dataset of invoices and payments following the characteristics of the actual data from an enterprise software system and mirroring many of the challenges described in C5, C7, and C8. Our dataset contains 15,521 invoices and 12,332 payments, and we experiment on 790 matching pairs and 1,210 pairs that do not match. We work with the data in two formats: (1) one format where each entity is represented by multiple tables and (2) another format where it is represented as one view that combines all tables, which is closer to the academic setting. Moreover, we formulate the entity matching task as a binary classification between rows similar to existing literature (Narayan et al., 2022; Peeters and Bizer, 2023). As such, we prompt the LLM to decide if two table rows (one payment and one invoice) match. For this experiment, we use a few-shot approach (as zero-shot performance was really low for enterprise data) and include one positive and one negative example (a match and a non-match) in the prompt. More examples did not improve the accuracy significantly. Afterward, we make the task incrementally more complex.
| Clean | + Errors | + Multi- Matches | + Multiple Tables | |
| GPT-4o-Mini | 0.98 | |||
| GPT-4o | 0.97 | |||
| Claude 3.5 Sonnet | 0.97 | 0.86 | ||
| Llama 3.1 Instruct | 0.99 | 0.95 | 0.72 |
Exp. 4: Increasing task complexity. In the following, we explain the different scenarios with results provided in Table 5:
1. (clean) First, we run the models on clean 1:1 matches using the single table views, which is closest to the academic setting. In this setting, the payment’s memo line includes the correct reference numbers, the payment amount is exactly as stated in the invoice, and the customer name is also identical. As a result, the models reach very high F1 scores in this setup, which we can see in Table 5 (Clean). Next, we incrementally increase the difficulty of matching payments to invoices based on the challenges we observe in enterprise data.
2. (+ errors) As mentioned before in C8, data errors in enterprises are different from academic datasets, making the matching harder. To show this, we add representative errors and discrepancies from real data to the source and target data that typically occur in real-world customer data, such as missing or additional digits in the reference numbers or (minor) discrepancies in the paid amount. As shown in Table 5 (+ Errors), these seemingly small inconsistencies already cause a noticeable drop in accuracy, highlighting the sensitivity of the models to such errors. We further analyze this case below in Exp. 5, which showcases that LLMs have real problems understanding the semantics of enterprise data.
3. (+ multi-matches) Next, we make the matching even harder and focus on multi-match cases as explained in C7, where one payment pays multiple invoices or multiple payments together pay one invoice, which is a setting that is already closer to real-world scenarios found in enterprises. As shown in Table 5 (+ Multi-Matches), this vastly increases the task’s difficulty. To understand the root cause of this drop in F1 scores, we further investigate the precision and recall to understand the high deviation in F1 scores between the models (see Figure 6). We find that while all models achieve very high precision, differences in F1 scores are primarily driven by variations in recall. This indicates that all of the models take a rather cautious approach, only predicting matches when they are highly certain. As a result, they prioritize high precision in matching and therefore miss many correct matches. While this high precision is sometimes preferred over high recall but low precision, the high rate of missed matches causes high manual efforts for enterprises to find the missing matches.
4. (+ multiple tables) Finally, to show the effect that business entities often span multiple tables as mentioned in C5, we represent invoices using multiple tables (as in the original SAP schema) instead of the single flat table view. While metadata about each invoice document is stored in one table, specific information like the amount and due date are stored in a second table, and information about customers is stored in another separate customer table. In this scenario, which now represents the actual enterprise task in its entire complexity, all LLMs see large performance decreases compared to the previous experiments, with the lowest performance in GPT-4o-Mini, indicating that the models have difficulties working with the complex data structures often used in enterprises.
Exp. 5: Data errors amplify task complexity? Our second experiment analyzes the impact of different error types on complex enterprise tasks like payment-to-invoice matching, showcasing some interesting findings which relate the task to typical data challenges in enterprise scenarios. In this experiment, we insert individual classes of errors into the clean data of Exp. 4 to study their effects. We focus on four typical cases: (1) we deduct up to USD 0.1 from the amount paid, (2) we remove or change digits in the assignment or (3) in the billing number, and (4) we slightly vary the partner name, for example with abbreviations like KL Technologies instead of Kim & Lee Technologies. The results for the different error classes shown in Figure 7 indicate that even minor discrepancies lead to performance degradation of the LLMs. Interestingly, errors in numerical fields only cause a relatively modest performance drop (around 5%), while discrepancies in business partner names result in a much more severe decline. This suggests that for enterprise data, the LLMs rely heavily on the availability of non-erroneous textual fields for matching, highlighting their inability to understand numerical and symbolic data, which is quite prominent in enterprise data as discussed in Section 4.1.
5.3. Experiments & Results (Case Study 2)
To investigate the impact of the compound tasks stated in C6, we now conduct several experiments based on our second case study (database integration) introduced at the beginning of this section.
Setup.
Similar to before, we are using a dataset based on real-world data for data integration from two companies. Company A’s database consists of a single table with 15 columns, selected as a subset of the SAP general customer master data table KNA1 with columns such as KUNNR, NAME1, and ERDAT. Company B’s database consists of two tables with a foreign key relationship, where the tables have self-descriptive column names such as Organization Name, Address, and Tax Number. We experiment with datasets of different sizes, from 50 to 300 customers, to analyze the effects of scaling the data. The customers in the two databases have an overlap of 60% (i.e., 40% of the customers exist only in one of the databases). To evaluate the predicted result table, we compare it with the ground truth result table of integrated customers and compute the cell-level accuracy (i.e. the fraction of correct cells).
Exp. 6: Error propagation in compound tasks. In our first experiment for this case study, we aim to analyze the effect of error propagation in compound tasks with multiple steps. For this, we formulate the task of customer data integration as a sequence of the following three steps, visualized in Figure 2 (center):
1. (schema matching) To integrate the data, we first perform schema matching to find the mapping between the columns of Company B’s database to the columns in Company A’s database. As in existing research, we formulate the task as a binary classification (Narayan et al., 2022; Parciak et al., 2024). Challenges arise because, for example, the address is represented in a single column in Company B, but is split into multiple fields in Company A. We measure the accuracy of this step as the fraction of correctly found matches for the columns in Company A since the output of the schema matching step is used to transform Company B’s database into Company A’s schema.
2. (entity matching) To merge the data correctly, we perform entity matching to find out which customers have entries in both companies’ databases. Similar to the existing literature (Narayan et al., 2022; Peeters and Bizer, 2023), we formulate entity matching as a binary classification task: Given a pair of rows, one from Company A’s table and one from Company B’s transformed table, the LLM determines whether they refer to the same customer.
3. (data integration) For the actual data integration, the final customer table is created by using the LLM to automatically merge the duplicate records found in both databases into a single row and transform Company B’s records into the format of Company A.
To evaluate the impact of error propagation when chaining the tasks, we run each step twice: once based on the output of the previous pipeline step, and once in a standalone setting; i.e., with the correct results (without errors) of the previous step as input. The results of this experiment are shown in Figure 8, where the results with error propagation (called Pipeline) are shown as colored bars, and the results of the Standalone setting (without error propagation) are shown as stacked, colorless bars.
For entity matching and data integration (the second and third steps in the pipeline), we observe that executing these tasks as a pipeline leads to a drop in accuracy compared to running them as standalone tasks. However, the decrease depends on the model. While for GPT-4o-Mini (smallest model), we see a severe drop due to error propagation, Claude and Llama see a much smaller drop.
Interestingly, the performance between the entity matching and data integration step in the pipeline does not decline and actually slightly improves for some models. This shows a surprising effect for LLMs: While the models make errors that propagate through the pipeline, they can, in some cases, correct earlier mistakes. For instance, some models fail to correctly associate the PSTLZ field with the postal code during schema matching. However, when presented with three sample rows from Company A during data integration, they are able to correctly populate the schema of Company A with the addresses stored in Company B’s data.
Overall, these results suggest that compound tasks clearly lead to a drop in accuracy due to error propagation, but the significance depends on the LLM. In some cases, the effects are even negligible.
Exp. 7: End-to-end task formulation. A high overhead in enterprise tasks stems from the fact that complex tasks must first be manually decomposed into individual steps by humans. Recent LLMs, however, can do reasoning to solve multi-step problems on their own. As such, one might ask the question if the manual decomposition of the problem is even needed, or if recent LLMs that are able to reason can break down the problems on their own?
To show evaluate this, we let the LLMs directly integrate the two customer databases with a single prompting step and compare the results to the pipelined task formulation from Exp. 6. We experiment with two prompting strategies: (1) w/o Steps provides a textual description with instructions on how to perform the integration but does not mention how to break down the task, whereas (2) with Steps explicitly mentions the necessary intermediate steps (schema matching, entity matching, and data integration) in the prompt to provide a hint towards how to execute them.
The results are shown in Figure 3. In this experiment, we include OpenAI’s o1 as an additional LLM with reasoning capabilities. We observe that most models perform better in the pipelined task formulation. Interestingly, only GPT-4o-Mini achieves better results end-to-end when compared to the pipelined execution of the experiment before. When comparing end-to-end w/o Steps and with Steps, we see mixed effects. While some LLMs benefit from describing the steps in the prompts, others (e.g., o1 and Claude) do not.
Exp. 8: End-to-end scaling. While the end-to-end formulation of compound tasks is attractive as it does not require manual task decomposition, it necessitates that the full tables are handed in as input to the LLM because some steps (e.g., a matching step between tables) might need access to all tuples.
As a last experiment in our customer integration case study, we investigate the effects of scaling the number of customers in the company databases. However, in the end-to-end task formulation, this leads to longer prompts that quickly exceed the available context windows. Among the models evaluated in this work, only OpenAI’s o1 reasoning model provides a sufficiently high output token limit (100K tokens), which is necessary to integrate more than 100 customers in an end-to-end manner.
Figure 10 presents the results of scaling the total number of customers from 50 up to 300. First, we observe a drastic decline in performance as the number of customers increases, suggesting that the model struggles with long-range dependencies and finding matches in large datasets. Moreover, it often fails to merge duplicate customers. When using 300 customers, the model even responds with ”Sorry, but I can’t fulfill that”. Overall, the ability of LLMs to handle enterprise-scale tasks end-to-end remains limited. In contrast, a formulation that breaks up compound tasks into individual steps, as in our previous experiment, can, in principle, scale to larger tables as tuples are handed in one by one (e.g., comparing all customers of databases A and B individually for matching).
5.4. Discussion
Learnings.
In our experiments, we experienced that automating enterprise tasks end-to-end with LLMs is extremely hard: First, we still need to rely on human effort as end-to-end approaches with LLMs have been found to be very brittle and not scalable. This is consistent with Ashury-Tahan et al. (2025), who found that even strong models often fail to perform robustly on complex tabular data tasks. In addition, we encountered problems in the output of some task formulations, like end-to-end, where the output is a table with multiple rows, since models often generate varying numbers of columns for individual rows. This typically does not happen with task formulations where the output is just a single row.
Second, much of the complexity of enterprise tasks is inherent to business needs and cannot be avoided, such as the need to deal with multi-match cases in entity-matching. Furthermore, we found that reducing complexity by working with database views improved the results. However, constructing such a view causes additional manual overheads as business objects in enterprises can be composed of tens to even hundreds of tables. Finally, some challenges like error propagation are interestingly not as severe as initially thought. Overall, however, given the current abilities, LLMs are still not sufficient to achieve the level of performance required for enterprise-scale data engineering.
Outlook.
To better handle the complexity of enterprise tasks, we think that LLMs on their own will not be sufficient since much of the complexity cannot be simplified, as discussed earlier. Instead, complex tasks need to be approached with systems that combine LLMs with other methods while also designing effective human-LLM interactions. Wornow et al. (2024) propose demonstrating workflows to foundation models, which is promising for automating frequently occurring tasks. However, such solutions still require manual overheads. Another promising direction is leveraging LLMs to first create structured plans for approaching compound tasks similar to recent advancements in using LLMs for query planning (Urban and Binnig, 2024; Liu et al., 2025) and then execute these plans step-wise using existing approaches (e.g., heuristics for schema matching) or even the LLM itself.
6. The Knowledge Challenge
Since LLMs are primarily pre-trained on public data, they lack information about a company’s internal business processes and policies, as well as about its proprietary tools and systems (Kayali et al., 2025). In this section, we analyze how this lack of company-specific knowledge in LLMs affects their accuracy on enterprise data engineering tasks.
6.1. Knowledge Challenges
We identify two main categories of challenges: (1) the lack of knowledge that is not available to LLMs but does exist in documentation or other company-internal sources, and (2) company-specific extensions of databases that are typically not documented at all.
C9: Proprietary but available knowledge. Enterprises often use proprietary tools and systems to address data engineering challenges. In contrast to well-established technologies like SQL, there is typically much less documentation about them available on the web. Therefore, it is reasonable to assume that LLMs trained on public data have little parametric knowledge about these tools and systems. One example is the domain-specific query language SIGNAL (Kampik et al., 2023), which SAP provides to its customers for exploring data about business processes. SIGNAL resembles SQL in many aspects but also includes domain-specific features and syntax differences tailored towards process mining. While the language itself is well-specified on its public documentation page, it has a much smaller user base than SQL and, therefore, a smaller online footprint regarding help pages, Q&A threads, and blog posts, resulting in limited understanding. Nevertheless, SIGNAL is frequently used by data engineers and must thus be well-supported by LLMs, for example, to translate natural language requests into SIGNAL queries.
C10: Proprietary and unavailable knowledge. Enterprise systems like those from SAP support company-specific changes and extensions to customize the system for individual customers. For example, customers can use hooks to add custom business logic. On the data level, customization means extending the database schema by adding customer-specific columns to existing tables or even additional tables to the customer namespace. In our analysis of real-world systems from SAP, we have come across thousands of such customer-defined tables. Since these changes are highly company-specific, they are typically not documented publicly—if they are documented at all. Moreover, customers sometimes “misuse” existing attributes of the standard SAP schema for different purposes. Such digressions from the public documentation pose significant challenges for data engineering with LLMs as well.
6.2. Experiments & Results
To analyze how the challenges C9-C10 affect data engineering with LLMs, we perform two experiments. (1) First, we apply LLMs to the task of translating natural language requests into the proprietary query language SIGNAL. As mentioned in C9, the specifics of SIGNAL provide a good example for proprietary knowledge that is available primarily in company-internal documentation. (2) In a second experiment, we exercise the case where no knowledge is available to enrich the context of LLMs, not even in internal documentation (C10). In this experiment, we want to determine to what extent LLMs understand customer-specific extensions to the standard SAP schema. With a simple probing task, we demonstrate how LLMs struggle with customizations and highlight their strong bias towards publicly documented parts of the schema.
Exp. 9: Text-to-SIGNAL. In our first experiment, we use LLMs to translate natural language requests into the proprietary query language SIGNAL, as shown in Figure 2 (right).
Setup.
For this experiment, we use a set of 200 randomly-sampled pairs of natural language requests and SIGNAL queries from a larger dataset in use at SAP. Given a natural language request, like Retrieve the number of unique invoices from ’defaultview-255’, we prompt the LLMs to generate a corresponding SIGNAL query. We compare three different approaches: First, we prompt the LLMs only with a short instruction explaining the task (Zero-shot). Next, we add three fixed representative example queries that we hand-picked from the SIGNAL documentation (+ Examples). Finally, we include both the examples as well as the full documentation of the SIGNAL language in the prompt (+ Documentation). For all three scenarios, we manually tune our prompts to alleviate obvious mistakes and hint at the differences between SIGNAL and SQL. To evaluate the generated queries, we compute the execution accuracy by executing the ground truth and predicted query and comparing their results.
Results.
Figure 11 shows the results of our Text-to-SIGNAL experiment (left bar groups). We can clearly see that only a small fraction (up to ) of the generated SIGNAL queries is correct. Moreover, we find that adding example queries to the prompt brings the biggest improvements in accuracy. By contrast, additionally including the full documentation of the SIGNAL language in the prompt improves the accuracy only for some models, indicating that simply providing the required documentation during inference is not a viable solution. Finally, a comparison with the results on popular Text-to-SQL benchmarks (Yu et al., 2018; Li et al., 2023) (Figure 11 right) shows that the accuracy for Text-to-SIGNAL is substantially lower.
Why are examples and documentation not helping? To better understand why exactly the models fail, we conduct an error analysis that categorizes the queries into those that execute with (1) correct or (2) incorrect results, and those that fail to execute because of (3) syntactic or (4) semantic errors. We focus on GPT-4o, the only model that could correctly translate a non-negligible number of requests in all three configurations. Figure 12 (left) shows that most of the generated queries fail to execute because of syntactic or semantic errors, with the largest number of failures caused by syntax errors. Adding example queries and documentation to the prompt causes the number of syntax errors to decrease. Nevertheless, we also see that the number of queries that contain semantic errors or return incorrect results increases as we add examples and documentation. A closer inspection of the generated queries reveals that the models often confuse parts of the documentation, like the column names used in explanations, with the request they have to translate. As a result, the number of fully correct queries is far from satisfactory even when including representative examples and the full SIGNAL documentation in the prompt.
A bias towards SQL. Another interesting effect we observe is that the models tend to generate SQL-specific syntax and ignore the specifics of SIGNAL. To further dive into the root causes, we manually analyze the syntax errors generated by GPT-4o in the (+ Documentation) setting and group them into three categories: (1) incorrect use of GROUP BY and ORDER BY, which in SIGNAL requires numerical indices instead of column names (unlike in SQL), (2) use of invalid characters, like an asterisk (*) in a count statement which exists in SQL but not in SIGNAL, and (3) incorrect structure of the overall statement. As shown in Figure 12 (right), we find that even when explicitly including rules in the prompt to avoid such SQL-related syntax errors, the models still generate erroneous SIGNAL queries. This behavior indicates a strong bias of LLMs towards SQL, which may be caused by its much higher prominence in their pre-training data. Similar problems could also exist in other scenarios where enterprise data or tasks differ only slightly from public knowledge, like currency conversion or complex data types.
Exp. 10: Schema customizations. In our next experiment, we analyze how LLMs perform when a task requires company-specific knowledge for which there is little to no explicit documentation.
Setup.
For this experiment, we use a scenario with customizations of the standard SAP schema and analyze to which extent the LLMs’ internal parametric knowledge already captures this knowledge. To see to which extent the parametric knowledge of LLMs is able to understand customized tables without introducing further task-specific difficulties, we choose a simple task setup: Given an SAP table name, we instruct the model to generate a list of all columns that belong to that table. That way, we can see to which extent LLMs know the table columns. In our analysis, we differentiate between columns that are part of the standard SAP schema and those that were defined by the customer. Our corpus consists of 2,000 tables from a real-world SAP system, of which 1,000 are part of the standard SAP schema and 1,000 were defined by the customer. We evaluate the predictions by computing F1 scores that measure the overlap between the predicted and true columns of each table.
Results.
Figure 13 shows the results broken down by the frequency of how often columns appear in the schema. As expected, the parametric knowledge about customer-defined columns is severely limited to certain columns like MANDT that must appear in most customer-defined tables. In contrast, the LLMs achieve higher F1 scores for columns from standard SAP tables. Still, we see that their parametric knowledge is strongly biased towards columns that appear in many tables and are thus more prominent in public documentation. Importantly, our introductory experiment (Figure 1 right) shows that this lack of enterprise-specific knowledge about customized columns severely impacts performance on downstream tasks like column type annotation.
6.3. Discussion
Learnings.
Our experiments show that the lack of enterprise knowledge affects the performance of LLMs. Even when including hand-picked examples or the full documentation in the prompt, LLMs are unable to actualize this information to solve enterprise tasks like Text-to-SIGNAL. An interesting finding here is that biases in the LLMs’ knowledge towards public sources, like their deep understanding of SQL, are hard to overcome even with explicit instructions, making LLMs difficult to adapt to a wide variety of specialized enterprise use cases without more heavy-weight solutions like fine-tuning that cause additional overheads.
Outlook.
As discussed before, our experiments show that simple out-of-the-box approaches like few-shot prompting and including relevant documentation in the prompt are insufficient. While we believe that more involved approaches, like fine-tuning models for specific enterprise use cases, could help them acquire the necessary domain knowledge, they require extensive work, for example to label the necessary data for fine-tuning. As such, we believe that future research should explore the direction of self-supervised pre-training of foundation models on large enterprise corpora, specifically on enterprise data and documentation. This could allow them to acquire a broad base of enterprise knowledge that enables them to solve complex enterprise tasks out of the box, which is an idea that has very recently gained some attraction in industry.444https://www.databricks.com/company/newsroom/press-releases/databricks-and-anthropic-sign-landmark-deal-bring-claude-models
7. Costs At Enterprise Scale
The cost of using LLMs at enterprise scale is an important orthogonal dimension to the challenges discussed before. In this section, we want to briefly highlight three main cost drivers: the large scale of the data, the algorithmic complexity of the tasks, and the high tokenizer fertility (Rust et al., 2021), which is a general problem on tabular data.
Data scale.
As described in Section 4, enterprise databases are substantially larger than typical data engineering benchmark datasets, with SAP databases often reaching multiple terabytes in size. Table 6 shows how this large size translates into high LLM costs. Processing one gigabyte of tabular data from our SAPCTA corpus already costs about USD 71. Note that this cost does not include generating any output tokens, which are substantially more expensive than input tokens, as shown in Table 1. Nevertheless, looking only at the input token costs already provides a rough estimate of the expected high costs. For example, processing the entire customer database with GPT-4o-Mini (OpenAI’s cheapest model) would cost USD 463T, and using OpenAI’s o1 model would cost USD 46M, as shown in Table 6. These high costs clearly prohibit using LLMs on large fractions of the enterprise data. This is especially problematic as our experiments in Sections 4-6 have shown that cheaper models like GPT-4o-Mini often underperform more expensive models.
Algorithmic complexity.
As a second cost driver, we want to highlight the algorithmic complexity of the ways in which many data engineering tasks are currently approached with LLMs. One example is the task of entity matching, which current approaches address by using LLMs to compare pairs of entities (Peeters and Bizer, 2023; Li et al., 2024). This approach has an algorithmic complexity of , where and are the numbers of entities in each table, making it intractable even for medium-sized tables. For example, matching just 1,000 payments to 1,000 invoices already costs USD 1,462, and matching 10,000 payments to 10,000 invoices would cost USD 146,165. Future work should therefore put a strong emphasis on reducing the costs of LLM-based approaches, for example by exploring task formulations with lower complexities (Wang et al., 2025), using LLMs together with smaller fine-tuned models (Liu et al., 2024b; Zhang et al., 2024b), or combining them with conventional approaches such as blocking (Zhang et al., 2020b).
| Tokens per Byte | USD per GB | |||
| GPT-4o-Mini | GPT-4o | o1 | ||
| Wikipedia | 0.23 | 34 | 574 | 3,442 |
| SAPCTA (CSV) | 0.47 | 71 | 1,181 | 7,085 |
| USD for entire database: | 462,923 | 7,715,380 | 46,292,282 | |
Pricing by OpenAI in Feb 2025.
Tokenizer fertility.
A final important cost driver is the high fertility of existing tokenizers on tabular data. The tokenizer fertility measures the number of tokens required to represent a single piece of text (Rust et al., 2021). Since tokenizers are trained to represent natural language text, their vocabulary contains tokens even for long words. By contrast, tabular enterprise data contains lots of symbols and numerical values. As shown in Table 6, this results in twice the number of tokens required to represent each byte of enterprise data compared to natural language text from Wikipedia, leading to twice the cost per gigabyte. Even if the cost of new models continues to decrease, the costs for processing tabular data will thus remain twice as high. Therefore, future work should focus on creating new tokenizers specifically designed for enterprise data, as well as on creating more efficient textual representations for such data (Trummer, 2024).
8. Conclusion
In this paper, we systematically analyzed the challenges involved in applying LLMs to real-world enterprise data engineering scenarios. Through experiments on a diverse set of tasks, we have shown how their effectiveness is constrained by enterprise-specific complexities along the dimensions of data, tasks, and knowledge, significantly impacting the reliability of LLM-driven automation in enterprises. We hope that our insights and learnings provide a helpful guide and inspiration for future efforts to make LLMs viable for enterprise data engineering and support their adoption in industry.
Acknowledgements.
This work has been supported by the BMBF and the state of Hesse as part of the NHR Program and the HMWK cluster project 3AI. It was also partially funded by the LOEWE Spitzenprofessur of the state of Hesse. We also thank DFKI Darmstadt, SAP, and hessian.AI as well as Dr. Alexey Streltsov, Matthias Urban, and Furkan Karakocaoglu for their support.References
- (1)
- Anthropic (2024) Anthropic. 2024. Claude 3.5 Sonnet Model Card Addendum.
- Ashury-Tahan et al. (2025) Shir Ashury-Tahan, Yifan Mai, Rajmohan C, et al. 2025. The Mighty ToRR: A Benchmark for Table Reasoning and Robustness. https://doi.org/10.48550/arXiv.2502.19412 arXiv:2502.19412 [cs].
- Barke et al. (2024) Shraddha Barke, Christian Poelitz, Carina Negreanu, et al. 2024. Solving Data-centric Tasks using Large Language Models. In Findings of the Association for Computational Linguistics: NAACL 2024, Kevin Duh, Helena Gomez, and Steven Bethard (Eds.). Association for Computational Linguistics, Mexico City, Mexico, 626–638. https://doi.org/10.18653/v1/2024.findings-naacl.41
- Bhagavatula et al. (2015) Chandra Sekhar Bhagavatula, Thanapon Noraset, and Doug Downey. 2015. TabEL: Entity Linking in Web Tables. In The Semantic Web - ISWC 2015. Springer International Publishing, Cham, 425–441. https://doi.org/10.1007/978-3-319-25007-6_25
- Bodensohn et al. (2024a) Jan-Micha Bodensohn, Ulf Brackmann, Liane Vogel, Anupam Sanghi, and Carsten Binnig. 2024a. Automating Enterprise Data Engineering with LLMs. In NeurIPS 2024 Third Table Representation Learning Workshop. https://openreview.net/forum?id=m85fYEJtDc
- Bodensohn et al. (2024b) Jan-Micha Bodensohn, Ulf Brackmann, Liane Vogel, Matthias Urban, Anupam Sanghi, and Carsten Binnig. 2024b. LLMs for Data Engineering on Enterprise Data. In Proceedings of Workshops at the 50th International Conference on Very Large Data Bases, VLDB 2024, Guangzhou, China, August 26-30, 2024. VLDB.org. https://vldb.org/workshops/2024/proceedings/TaDA/TaDA.4.pdf
- Brown et al. (2020) Tom Brown, Benjamin Mann, Nick Ryder, et al. 2020. Language Models are Few-Shot Learners. In Advances in Neural Information Processing Systems, H. Larochelle, M. Ranzato, R. Hadsell, M. F. Balcan, and H. Lin (Eds.), Vol. 33. Curran Associates, Inc., 1877–1901. https://proceedings.neurips.cc/paper_files/paper/2020/file/1457c0d6bfcb4967418bfb8ac142f64a-Paper.pdf
- Chen et al. (2024) Peter Baile Chen, Fabian Wenz, Yi Zhang, et al. 2024. BEAVER: An Enterprise Benchmark for Text-to-SQL. https://doi.org/10.48550/arXiv.2409.02038 arXiv:2409.02038 [cs].
- Dao et al. (2022) Tri Dao, Dan Fu, Stefano Ermon, Atri Rudra, and Christopher Ré. 2022. FlashAttention: Fast and Memory-Efficient Exact Attention with IO-Awareness. In Advances in Neural Information Processing Systems, S. Koyejo, S. Mohamed, A. Agarwal, D. Belgrave, K. Cho, and A. Oh (Eds.), Vol. 35. Curran Associates, Inc., 16344–16359. https://proceedings.neurips.cc/paper_files/paper/2022/file/67d57c32e20fd0a7a302cb81d36e40d5-Paper-Conference.pdf
- DeepSeek-AI et al. (2025) DeepSeek-AI, Daya Guo, Dejian Yang, et al. 2025. DeepSeek-R1: Incentivizing Reasoning Capability in LLMs via Reinforcement Learning. arXiv:2501.12948 [cs.CL] https://arxiv.org/abs/2501.12948
- Deng et al. (2020) Xiang Deng, Huan Sun, Alyssa Lees, You Wu, and Cong Yu. 2020. TURL: table understanding through representation learning. Proc. VLDB Endow. 14, 3 (Nov. 2020), 307–319. https://doi.org/10.14778/3430915.3430921
- Eisenschlos et al. (2021) Julian Eisenschlos, Maharshi Gor, Thomas Müller, and William Cohen. 2021. MATE: Multi-view Attention for Table Transformer Efficiency. In Proceedings of the 2021 Conference on Empirical Methods in Natural Language Processing. Association for Computational Linguistics, Online and Punta Cana, Dominican Republic, 7606–7619. https://doi.org/10.18653/v1/2021.emnlp-main.600
- Frieder et al. (2023) Simon Frieder, Luca Pinchetti, Chevalier Chevalier, et al. 2023. Mathematical Capabilities of ChatGPT. In Advances in Neural Information Processing Systems, A. Oh, T. Naumann, A. Globerson, K. Saenko, M. Hardt, and S. Levine (Eds.), Vol. 36. Curran Associates, Inc., 27699–27744. https://proceedings.neurips.cc/paper_files/paper/2023/file/58168e8a92994655d6da3939e7cc0918-Paper-Datasets_and_Benchmarks.pdf
- Gao et al. (2024) Dawei Gao, Haibin Wang, Yaliang Li, et al. 2024. Text-to-SQL Empowered by Large Language Models: A Benchmark Evaluation. Proc. VLDB Endow. 17, 5 (Jan. 2024), 1132–1145. https://doi.org/10.14778/3641204.3641221
- Gao et al. (2025) Yingqi Gao, Yifu Liu, Xiaoxia Li, et al. 2025. A Preview of XiYan-SQL: A Multi-Generator Ensemble Framework for Text-to-SQL. https://doi.org/10.48550/arXiv.2411.08599 arXiv:2411.08599 [cs].
- Grattafiori et al. (2024) Aaron Grattafiori, Abhimanyu Dubey, Abhinav Jauhri, et al. 2024. The Llama 3 Herd of Models. https://doi.org/10.48550/arXiv.2407.21783 arXiv:2407.21783 [cs].
- He et al. (2025) Xinyi He, Yihao Liu, Mengyu Zhou, et al. 2025. TableLoRA: Low-rank Adaptation on Table Structure Understanding for Large Language Models. https://doi.org/10.48550/arXiv.2503.04396 arXiv:2503.04396 [cs].
- Herzig et al. (2020) Jonathan Herzig, Pawel Krzysztof Nowak, Thomas Müller, Francesco Piccinno, and Julian Eisenschlos. 2020. TaPas: Weakly Supervised Table Parsing via Pre-training. In Proceedings of the 58th Annual Meeting of the Association for Computational Linguistics. Association for Computational Linguistics, Online, 4320–4333. https://doi.org/10.18653/v1/2020.acl-main.398
- Hulsebos et al. (2021) Madelon Hulsebos, Cağatay Demiralp, and Paul Demiralp. 2021. GitTables benchmark - column type detection. https://doi.org/10.5281/zenodo.5706316
- Hulsebos et al. (2023) Madelon Hulsebos, Cagatay Demiralp, and Paul Groth. 2023. GitTables: A Large-Scale Corpus of Relational Tables. Proceedings of the ACM on Management of Data 1, 1 (May 2023), 30:1–30:17. https://doi.org/10.1145/3588710
- Hulsebos et al. (2019) Madelon Hulsebos, Kevin Hu, Michiel Bakker, et al. 2019. Sherlock: A Deep Learning Approach to Semantic Data Type Detection. In Proceedings of the 25th ACM SIGKDD International Conference on Knowledge Discovery & Data Mining (KDD ’19). Association for Computing Machinery, New York, NY, USA, 1500–1508. https://doi.org/10.1145/3292500.3330993
- Jaimovitch-López et al. (2022) Gonzalo Jaimovitch-López, Cèsar Ferri, José Hernández-Orallo, Fernando Martínez-Plumed, and María José Ramírez-Quintana. 2022. Can language models automate data wrangling? Machine Learning 112, 6 (2022), 2053–2082. https://doi.org/10.1007/s10994-022-06259-9
- Kampik et al. (2023) Timotheus Kampik, Andre Lücke, Jörn Horstmann, Mark Wheeler, and David Eickhoff. 2023. SIGNAL – The SAP Signavio Analytics Query Language. https://doi.org/10.48550/arXiv.2304.06811 arXiv:2304.06811 [cs].
- Kang and Naughton (2003) Jaewoo Kang and Jeffrey F. Naughton. 2003. On schema matching with opaque column names and data values. In Proceedings of the 2003 ACM SIGMOD international conference on Management of data (SIGMOD ’03). Association for Computing Machinery, New York, NY, USA, 205–216. https://doi.org/10.1145/872757.872783
- Kayali et al. (2024) Moe Kayali, Anton Lykov, Ilias Fountalis, Nikolaos Vasiloglou, Dan Olteanu, and Dan Suciu. 2024. Chorus: Foundation Models for Unified Data Discovery and Exploration. Proc. VLDB Endow. 17, 8 (April 2024), 2104–2114. https://doi.org/10.14778/3659437.3659461
- Kayali et al. (2025) Moe Kayali, Fabian Wenz, Nesime Tatbul, and Çağatay Demiralp. 2025. Mind the Data Gap: Bridging LLMs to Enterprise Data Integration. In 15th Annual Conference on Innovative Data Systems Research.
- Klein et al. (2024) Tassilo Klein, Clemens Biehl, Margarida Costa, Andre Sres, Jonas Kolk, and Johannes Hoffart. 2024. SALT: Sales Autocompletion Linked Business Tables Dataset. In NeurIPS 2024 Third Table Representation Learning Workshop. https://openreview.net/forum?id=UZbELpkWIr
- Konda et al. (2016) Pradap Konda, Sanjib Das, Paul Suganthan G. C., et al. 2016. Magellan: toward building entity matching management systems over data science stacks. Proceedings of the VLDB Endowment 9, 13 (Sept. 2016), 1581–1584. https://doi.org/10.14778/3007263.3007314
- Korini and Bizer (2023) Keti Korini and Christian Bizer. 2023. Column Type Annotation using ChatGPT. In Joint Proceedings of Workshops at the 49th International Conference on Very Large Data Bases (VLDB 2023), Vancouver, Canada, August 28 - September 1, 2023 (CEUR Workshop Proceedings), Vol. 3462. CEUR-WS.org. https://ceur-ws.org/Vol-3462/TADA1.pdf
- Korini et al. (2022) Keti Korini, Ralph Peeters, and Christian Bizer. 2022. SOTAB: The WDC Schema.org Table Annotation Benchmark. SemTab@ISWC 2022, October 23–27, 2022, Hangzhou, China (Virtual) 3320 (2022), 14–19.
- Langenecker et al. (2023) Sven Langenecker, Christoph Sturm, Christian Schalles, and Carsten Binnig. 2023. SportsTables: A new Corpus for Semantic Type Detection. (2023). https://doi.org/10.18420/BTW2023-68 ISBN: 9783885797258 Publisher: Gesellschaft für Informatik e.V.
- Li et al. (2023) Jinyang Li, Binyuan Hui, Ge Qu, et al. 2023. Can LLM Already Serve as A Database Interface? A BIg Bench for Large-Scale Database Grounded Text-to-SQLs. In Advances in Neural Information Processing Systems 36: Annual Conference on Neural Information Processing Systems 2023, NeurIPS 2023, New Orleans, LA, USA, December 10 - 16, 2023, Alice Oh, Tristan Naumann, Amir Globerson, Kate Saenko, Moritz Hardt, and Sergey Levine (Eds.). http://papers.nips.cc/paper_files/paper/2023/hash/83fc8fab1710363050bbd1d4b8cc0021-Abstract-Datasets_and_Benchmarks.html
- Li et al. (2024) Peng Li, Yeye He, Dror Yashar, et al. 2024. Table-GPT: Table Fine-tuned GPT for Diverse Table Tasks. Proc. ACM Manag. Data 2, 3 (May 2024), 176:1–176:28. https://doi.org/10.1145/3654979
- Liu et al. (2025) Chunwei Liu, Matthew Russo, Michael Cafarella, et al. 2025. Palimpzest: Optimizing AI-Powered Analytics with Declarative Query Processing. In 15th Annual Conference on Innovative Data Systems Research.
- Liu et al. (2024a) Nelson F. Liu, Kevin Lin, John Hewitt, et al. 2024a. Lost in the Middle: How Language Models Use Long Contexts. Trans. Assoc. Comput. Linguistics 12 (2024), 157–173. https://doi.org/10.1162/TACL_A_00638
- Liu et al. (2023) Pengfei Liu, Weizhe Yuan, Jinlan Fu, Zhengbao Jiang, Hiroaki Hayashi, and Graham Neubig. 2023. Pre-train, Prompt, and Predict: A Systematic Survey of Prompting Methods in Natural Language Processing. Comput. Surveys 55, 9 (Jan. 2023), 195:1–195:35. https://doi.org/10.1145/3560815
- Liu et al. (2024b) Yurong Liu, Eduardo Pena, Aecio Santos, Eden Wu, and Juliana Freire. 2024b. Magneto: Combining Small and Large Language Models for Schema Matching. https://doi.org/10.48550/arXiv.2412.08194 arXiv:2412.08194 [cs].
- Mei et al. (2021) Yinan Mei, Shaoxu Song, Chenguang Fang, Haifeng Yang, Jingyun Fang, and Jiang Long. 2021. Capturing Semantics for Imputation with Pre-trained Language Models. In 2021 IEEE 37th International Conference on Data Engineering (ICDE). IEEE, Chania, Greece, 61–72. https://doi.org/10.1109/ICDE51399.2021.00013
- Meta (2024) Meta. 2024. Llama 3.1 Model Card.
- Mudgal et al. (2018) Sidharth Mudgal, Han Li, Theodoros Rekatsinas, et al. 2018. Deep learning for entity matching: A design space exploration. In Proceedings of the 2018 international conference on management of data. 19–34.
- Narayan et al. (2022) Avanika Narayan, Ines Chami, Laurel Orr, and Christopher Ré. 2022. Can Foundation Models Wrangle Your Data? Proceedings of the VLDB Endowment 16, 4 (Dec. 2022), 738–746. https://doi.org/10.14778/3574245.3574258
- OpenAI (2024a) OpenAI. 2024a. GPT-4o System Card.
- OpenAI (2024b) OpenAI. 2024b. Learning to Reason with LLMs. https://openai.com/index/learning-to-reason-with-llms/.
- OpenAI (2024c) OpenAI. 2024c. OpenAI o1 System Card.
- OpenAI et al. (2024) OpenAI, Josh Achiam, Steven Adler, et al. 2024. GPT-4 Technical Report. http://arxiv.org/abs/2303.08774 arXiv:2303.08774 [cs].
- Papenbrock et al. (2015) Thorsten Papenbrock, Arvid Heise, and Felix Naumann. 2015. Progressive Duplicate Detection. IEEE Transactions on Knowledge and Data Engineering 27, 5 (May 2015), 1316–1329. https://doi.org/10.1109/TKDE.2014.2359666
- Parciak et al. (2024) Marcel Parciak, Brecht Vandevoort, Frank Neven, Liesbet M. Peeters, and Stijn Vansummeren. 2024. Schema Matching with Large Language Models: an Experimental Study. In Proceedings of Workshops at the 50th International Conference on Very Large Data Bases, VLDB 2024, Guangzhou, China, August 26-30, 2024. VLDB.org. https://vldb.org/workshops/2024/proceedings/TaDA/TaDA.8.pdf
- Peeters and Bizer (2023) Ralph Peeters and Christian Bizer. 2023. Using ChatGPT for Entity Matching. In New Trends in Database and Information Systems - ADBIS 2023 Short Papers, Doctoral Consortium and Workshops: AIDMA, DOING, K-Gals, MADEISD, PeRS, Barcelona, Spain, September 4-7, 2023, Proceedings (Communications in Computer and Information Science), Vol. 1850. Springer, 221–230. https://doi.org/10.1007/978-3-031-42941-5_20
- Peeters et al. (2024) Ralph Peeters, Reng Chiz Der, and Christian Bizer. 2024. WDC Products: A Multi-Dimensional Entity Matching Benchmark. In Proceedings 27th International Conference on Extending Database Technology, EDBT 2024, Paestum, Italy, March 25 - March 28, Letizia Tanca, Qiong Luo, Giuseppe Polese, Loredana Caruccio, Xavier Oriol, and Donatella Firmani (Eds.). OpenProceedings.org, 22–33. https://doi.org/10.48786/EDBT.2024.03
- Qu et al. (2025) Jingang Qu, David Holzmüller, Gaël Varoquaux, and Marine Le Morvan. 2025. TabICL: A Tabular Foundation Model for In-Context Learning on Large Data. https://doi.org/10.48550/arXiv.2502.05564 arXiv:2502.05564 [cs].
- Rust et al. (2021) Phillip Rust, Jonas Pfeiffer, Ivan Vulić, Sebastian Ruder, and Iryna Gurevych. 2021. How Good is Your Tokenizer? On the Monolingual Performance of Multilingual Language Models. In Proceedings of the 59th Annual Meeting of the Association for Computational Linguistics and the 11th International Joint Conference on Natural Language Processing (Volume 1: Long Papers), Chengqing Zong, Fei Xia, Wenjie Li, and Roberto Navigli (Eds.). Association for Computational Linguistics, Online, 3118–3135. https://doi.org/10.18653/v1/2021.acl-long.243
- Savelieva et al. (2022) Alexandra Savelieva, Andreas Mueller, Avrilia Floratou, et al. 2022. The Need for Tabular Representation Learning: An Industry Perspective. Table Representation Learning Workshop at NeurIPS 2022 (2022).
- Shazeer et al. (2017) Noam Shazeer, Azalia Mirhoseini, Krzysztof Maziarz, et al. 2017. Outrageously Large Neural Networks: The Sparsely-Gated Mixture-of-Experts Layer. In International Conference on Learning Representations.
- Singha et al. (2023) Ananya Singha, José Cambronero, and Sumit Gulwani. 2023. Tabular Representation, Noisy Operators, and Impacts on Table Structure Understanding Tasks in LLMs. Table Representation Learning Workshop at NeurIPS 2023 (2023).
- Sui et al. (2024) Yuan Sui, Mengyu Zhou, Mingjie Zhou, Shi Han, and Dongmei Zhang. 2024. Table Meets LLM: Can Large Language Models Understand Structured Table Data? A Benchmark and Empirical Study. In Proceedings of the 17th ACM International Conference on Web Search and Data Mining. ACM, Merida Mexico, 645–654. https://doi.org/10.1145/3616855.3635752
- Trummer (2024) Immanuel Trummer. 2024. Generating Succinct Descriptions of Database Schemata for Cost-Efficient Prompting of Large Language Models. Proc. VLDB Endow. 17, 11 (Aug. 2024), 3511–3523. https://doi.org/10.14778/3681954.3682017
- Urban and Binnig (2024) Matthias Urban and Carsten Binnig. 2024. CAESURA: Language Models as Multi-Modal Query Planners. In 14th Conference on Innovative Data Systems Research, CIDR 2024, Chaminade, HI, USA, January 14-17, 2024. www.cidrdb.org. https://www.cidrdb.org/cidr2024/papers/p14-urban.pdf
- Vogel et al. (2022) Liane Vogel, Benjamin Hilprecht, and Carsten Binnig. 2022. Towards Foundation Models for Relational Databases [Vision Paper]. In NeurIPS 2022 First Table Representation Workshop. https://openreview.net/forum?id=s1KlNOQq71_
- Vogelsgesang et al. (2018) Adrian Vogelsgesang, Michael Haubenschild, Jan Finis, et al. 2018. Get Real: How Benchmarks Fail to Represent the Real World. In Proceedings of the Workshop on Testing Database Systems. ACM, Houston TX USA, 1–6. https://doi.org/10.1145/3209950.3209952
- Wang et al. (2025) Tianshu Wang, Xiaoyang Chen, Hongyu Lin, et al. 2025. Match, Compare, or Select? An Investigation of Large Language Models for Entity Matching. In Proceedings of the 31st International Conference on Computational Linguistics, Owen Rambow, Leo Wanner, Marianna Apidianaki, Hend Al-Khalifa, Barbara Di Eugenio, and Steven Schockaert (Eds.). Association for Computational Linguistics, Abu Dhabi, UAE, 96–109. https://aclanthology.org/2025.coling-main.8/
- Wei et al. (2022) Jason Wei, Maarten Bosma, Vincent Y. Zhao, et al. 2022. Finetuned Language Models are Zero-Shot Learners. In The Tenth International Conference on Learning Representations, ICLR 2022, Virtual Event, April 25-29, 2022. OpenReview.net. https://openreview.net/forum?id=gEZrGCozdqR
- Wornow et al. (2024) Michael Wornow, Avanika Narayan, Krista Opsahl-Ong, Quinn McIntyre, Nigam Shah, and Christopher Ré. 2024. Automating the Enterprise with Foundation Models. Proc. VLDB Endow. 17, 11 (2024), 2805–2812. https://doi.org/10.14778/3681954.3681964
- Wu et al. (2025) Zhaomin Wu, Shida Wang, Ziyang Wang, and Bingsheng He. 2025. Learning Relational Tabular Data without Shared Features. https://doi.org/10.48550/arXiv.2502.10125 arXiv:2502.10125 [cs].
- Yang et al. (2022) Jingfeng Yang, Aditya Gupta, Shyam Upadhyay, Luheng He, Rahul Goel, and Shachi Paul. 2022. TableFormer: Robust Transformer Modeling for Table-Text Encoding. In Proceedings of the 60th Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers). Association for Computational Linguistics, Dublin, Ireland, 528–537. https://doi.org/10.18653/v1/2022.acl-long.40
- Yin et al. (2020) Pengcheng Yin, Graham Neubig, Wen-tau Yih, and Sebastian Riedel. 2020. TaBERT: Pretraining for Joint Understanding of Textual and Tabular Data. In Proceedings of the 58th Annual Meeting of the Association for Computational Linguistics. Association for Computational Linguistics, Online, 8413–8426. https://doi.org/10.18653/v1/2020.acl-main.745
- Yu et al. (2018) Tao Yu, Rui Zhang, Kai Yang, et al. 2018. Spider: A Large-Scale Human-Labeled Dataset for Complex and Cross-Domain Semantic Parsing and Text-to-SQL Task. In Proceedings of the 2018 Conference on Empirical Methods in Natural Language Processing, Brussels, Belgium, October 31 - November 4, 2018, Ellen Riloff, David Chiang, Julia Hockenmaier, and Jun’ichi Tsujii (Eds.). Association for Computational Linguistics, 3911–3921. https://doi.org/10.18653/V1/D18-1425
- Zhang et al. (2020a) Dan Zhang, Madelon Hulsebos, Yoshihiko Suhara, Çağatay Demiralp, Jinfeng Li, and Wang-Chiew Tan. 2020a. Sato: contextual semantic type detection in tables. Proceedings of the VLDB Endowment 13, 12 (Aug. 2020), 1835–1848. https://doi.org/10.14778/3407790.3407793
- Zhang et al. (2024a) Haochen Zhang, Yuyang Dong, Chuan Xiao, and Masafumi Oyamada. 2024a. Large Language Models as Data Preprocessors. In Proceedings of Workshops at the 50th International Conference on Very Large Data Bases, VLDB 2024, Guangzhou, China, August 26-30, 2024. VLDB.org. https://vldb.org/workshops/2024/proceedings/TaDA/TaDA.11.pdf
- Zhang et al. (2020b) Wei Zhang, Hao Wei, Bunyamin Sisman, Xin Luna Dong, Christos Faloutsos, and Davd Page. 2020b. AutoBlock: A Hands-off Blocking Framework for Entity Matching. In Proceedings of the 13th International Conference on Web Search and Data Mining (WSDM ’20). Association for Computing Machinery, New York, NY, USA, 744–752. https://doi.org/10.1145/3336191.3371813
- Zhang et al. (2024b) Zeyu Zhang, Paul Groth, Iacer Calixto, and Sebastian Schelter. 2024b. Directions Towards Efficient and Automated Data Wrangling with Large Language Models. In 2024 IEEE 40th International Conference on Data Engineering Workshops (ICDEW). 301–304. https://doi.org/10.1109/ICDEW61823.2024.00044 ISSN: 2473-3490.
- Zhu et al. (2019) Erkang Zhu, Dong Deng, Fatemeh Nargesian, and Renée J. Miller. 2019. JOSIE: Overlap Set Similarity Search for Finding Joinable Tables in Data Lakes. In Proceedings of the 2019 International Conference on Management of Data (SIGMOD ’19). Association for Computing Machinery, New York, NY, USA, 847–864. https://doi.org/10.1145/3299869.3300065